OpenStudy (anonymous):

Multiply and simplify: √2(√8+4√3) HELP ME!! >=/

OpenStudy (anonymous):

Is this your question:\[\sqrt{2(\sqrt{8}+4\sqrt{3})}\]

OpenStudy (anonymous):

no . i copied it exactly how its up

OpenStudy (anonymous):

do u get it?

OpenStudy (anonymous):

I think so the whole thing isn't squared correct? So like this: \[\sqrt{2}(\sqrt{8}+4\sqrt{3})\]

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

Okay you have to take into consideration that when you have two numbers that are raised to the same root that: \[\sqrt{a}*\sqrt{b}=\sqrt{ab}\]

OpenStudy (anonymous):

So do this help you answer your question?

OpenStudy (anonymous):

not really

OpenStudy (anonymous):

It mean that your are able to multiply the 2 number under the square root. Let me use real numbers as an example \[\sqrt{3}*\sqrt{3}=\sqrt{9}=3\] Does that make it clearer?

OpenStudy (anonymous):

yeha

OpenStudy (anonymous):

so thats the answer?

OpenStudy (anonymous):

Okay let me know what you get

OpenStudy (anonymous):

No it's not you have to use the same method and apply it to your question.

OpenStudy (anonymous):

4+4 square root of 6?

OpenStudy (anonymous):

yup :).